A BLOCK OF MASS 2 KG IS PLACED ON THE FLOOR. THE COEFFICIENT OF STATIC FRICTION IS 0.4. IF  A FORCE OF 1.25 ACTS FOR 8 SEC. FINAL VELOCITY OF THE BODY IS

Grade:9

1 Answers

Arun
6 years ago

We have, coefficient of static friction, µs = 0.4

Mass, m = 2 kg

Force applied, F = 2.5 N

Maximum static frictional force, fmax= µs × m × g = 0.4 × 2 × 9.8

=> fmax = 7.84 N

Since, the applied force is less than the maximum static frictional force, the frictional force on the block is equal to the applied force = 2.5 N. This is according to the fact that static friction is a self adjusting force.